OpenStudy (republican31):

Write an equation of the line with the given slope and y-intercept in standard form. slope: 0.5; y-intercept: 7.5 Is 0.5x + y = 7.5 correct or a start or anything?

OpenStudy (anonymous):

y=mx+b where m is slope If you get y by itself y=-0.5x+7.5

OpenStudy (republican31):

Thank you but how do I make it standard form? Standard form is Ax + By = C but I can't remember what a and b are. I think A is the slope and B might be the y intercept?

OpenStudy (anonymous):

A and B are any number.

OpenStudy (anonymous):

You got the y intercept right. But something is wrong. Can you see it?

OpenStudy (republican31):

Not really sorry. I'm a little confused on standard form.

OpenStudy (anonymous):

standard form is Ax+By=C you are right on that. Hint:y=mx+b is wrong. where m is SLOPE

OpenStudy (republican31):

In your example for slope intercept form, you have -0.5 for m.

OpenStudy (anonymous):

They do not want -0.5 for m right? They want 0.5

OpenStudy (republican31):

Ya.

OpenStudy (anonymous):

Problem says slope: 0.5 slope=m

OpenStudy (anonymous):

So let me see your new standard form

OpenStudy (republican31):

I don't understand. I said 0.5x + y = 7.5 above, I think.

OpenStudy (republican31):

I just don't understand how to convert slope intercept to standard. I can't remember it and I've looked at a few websites and it doesn't seem to show me how, just provide practice problems.

OpenStudy (anonymous):

0.5x+y=7.5 turns to y=-0.5x+7.5 if you make y by itself and y=mx+b where m is slope m=-0.5 m=slope

OpenStudy (republican31):

Oh okay I see. I didn't understand what you were going for there. Alright so what would I have to do now?

OpenStudy (anonymous):

Make m the right slope.

OpenStudy (republican31):

-0.5x + y = 7.5?

OpenStudy (anonymous):

Yes, that is correct. Nice job.

OpenStudy (republican31):

Alright thank you! So I just need to start the proper way with slope intercept then convert to standard is all?

OpenStudy (anonymous):

Yes, slope intercept first then standard. It makes everything easier at the beginning.

OpenStudy (republican31):

Alright thank you so much!

OpenStudy (anonymous):

I think you guy missed the final step. For standard form, all the numbers have to be integers, and according to most sources, the coefficient of x must be positive. So you need to multiply by -2 to get rid of the decimals and make the "a" positive

OpenStudy (republican31):

By -2? Isn't it usually be -1 to keep the numbers the same? @Shalante ?

OpenStudy (anonymous):

If the numbers were integers, yes you'd multiply by -1. But you have decimals. You need to multiply. If they were fractions they'd be